OpenStudy (anonymous):

Find the value of a for which the expression (2a+3)x^2-6x+4-a is a perfect square.

OpenStudy (jdoe0001):

do you know what a "perfect square trinomial" is?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

how about that

OpenStudy (jdoe0001):

ok... do you know what a perfect square is?

OpenStudy (jdoe0001):

http://www.youtube.com/watch?v=aM1qf2QN8YE

OpenStudy (anonymous):

couldn't get it

OpenStudy (anonymous):

please help solve

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

\[x ^{2}-\frac{ 6}{2a+3 }x+\frac{ 4-a }{2a+3}=0\] \[x ^{2}-\frac{ 6 }{2a+3 }x+\left( \frac{ 3 }{2a+3 } \right)^{2}=\frac{ a-4 }{2a+3}+\left( \frac{ 3 }{2a+3}\right)^{2} \] put R.H.S=0 and find the values of a.

OpenStudy (anonymous):

\[\left( x-\frac{ 3 }{2a+3 } \right)^{2}=\frac{ 2a ^{2}+3a-8a-12+9 }{\left( 2a+3 \right)^{2}}\] \[=\frac{ 2a ^{2}-5a-3 }{\left( 2a+3 \right)^{2} }\] put R.H.S.=0 and find the values of a.

Parth (parthkohli):

Another way to do it is discriminants (Coeff of x)^2 - [4 * (coeff of x^2) * (constant)] = 0 for it to be a perfect square

hero (hero):

I knew there was an easier way, just couldn't think of it